# Fairness and Bias in Algorithmic Hiring: a Multidisciplinary Survey

### v2

@@ Abstract.
copyright: acmcopyrightjournalyear: 2023doi: XXXXXXX.XXXXXXXconference: Make sure to enter the correct conference title from your rights confirmation emai; September 2023; price: 15.00isbn: 978-1-4503-XXXX-X/18/06ccs: General and reference Surveys and overviewsccs: Information systems Data management systemsccs: Social and professional topics Computing / technology policy
@@ 4.2. Notation
$x\in\mathcal{X}$
a vector of non-sensitive attributes
$s\in\mathcal{S}$
a vector of sensitive attributes
$s=g$
a sensitive attribute value
$s={g}^{\mathsf{c}}$
complement of a sensitive attribute value
$y\in\mathcal{Y}$
target variable from domain $\mathcal{Y}$
$i=(x_{i},s_{i},y_{i})$
a data point or item
$\hat{y}=f(x)$
a classifier $f:\mathcal{X}\rightarrow\mathcal{Y}$ issuing predictions in $\mathcal{Y}$ for data points in $\mathcal{X}$
$f_{\text{soft}}(x)$
a soft classifier, from which predictions $\hat{y}=f(x)$ can be derived through thresholding
$\tau=\text{argsort}(f_{\text{soft}}(x))$
ranking of items, typically sorted by target estimates
$i=\tau(k)$
item at rank $k$ in $\tau$
$N$
number of items
$N_{g}$
number of items in $g$
$N_{g}^{k}$
number of items in $g$ in the top $k$ positions of a ranking
$D_{g}$
desired representation for group $g$
$h(x)$
a classifier $h:\mathcal{X}\rightarrow\mathcal{S}$ issuing predictions in $\mathcal{S}$ for data points
$h_{\text{soft}}(x)$
a soft classifier for sensitive attributes
@@ 4.3.2. Accuracy fairness
(Geyik et al., 2019)
outcome No No D
$\operatorname{NDKL}$
(Geyik et al., 2019; Arafan et al., 2022)
outcome No
$\operatorname{DI}$
(Booth et al., 2021; Burke et al., 2021; Köchling et al., 2021, 2021; Wilson et al., 2021; Deshpande et al., 2020)
outcome No
$\operatorname{DD}$
(Rus et al., 2022)
outcome No No
$\operatorname{RPP}$
(Ali et al., 2019; Imana et al., 2021; Peña et al., 2020)
outcome No No No
$\operatorname{DRD}$
(Zhang et al., 2022)
outcome No
$\operatorname{SKL}$
(Peña et al., 2020)
outcome No No
$\operatorname{LRR}$
(Chen et al., 2018a)
outcome $x$ No No D
$\operatorname{TPRD}$
(De-Arteaga et al., 2019; Hemamou and Coleman, 2022)
outcome $y$ No No
$\operatorname{FNRR}$
(Köchling et al., 2021)
outcome $y$ No No
$\operatorname{MED}$
(Singhania et al., 2020)
outcome $y$
$\operatorname{RMS}$
(Hemamou and Coleman, 2022)
outcome $y$ No No No
$\operatorname{xEO}$
(Nandy et al., 2022)
outcome $y$ No No No
$\operatorname{MAE}$
(Yan et al., 2020; Singhania et al., 2020)
accuracy $y$ No No
$\operatorname{BCRD}$
(Köchling et al., 2021)
accuracy $y$ No No No
$\operatorname{MID}$
(Köchling et al., 2021)
accuracy $y$ No No No
$\operatorname{SD}$
(Rus et al., 2022)
impact No No
$\operatorname{GBS}$
(Hu et al., 2022)
representational No n.a No
$\operatorname{sAUC}$
(Parasurama et al., 2022; Parasurama and Sedoc, 2021; Booth et al., 2021; Hemamou et al., 2021; Rus et al., 2022; Hemamou and Coleman, 2022)
process No n.a. No n.a.
$\operatorname{GTR}$
(Parasurama et al., 2022)
process $x$ No n.a. No n.a.
$\operatorname{MIA}$
(Yan et al., 2020)
process $y$ No n.a.
@@ 4.4. Discussion
Assorted fairness flavors. Most of the fairness measures considered in hiring focus on outcome equity; the field is strongly influenced by the 80% rule (Section 8.2), which is appreciated as a quantitative rule of thumb, but also often criticized (Wax, 2011; EEOC - US Equal Employment Opportunity Commission, 2023). Indeed, outcome fairness is based on a narrow view of algorithmic hiring as a single decision point abstracted away from its context, where equity is purely a function of algorithmic estimates and (sometimes) their similarity to target variables approximating a ground truth. Similar criticism can be moved to accuracy fairness, with the additional limitation that it is less interpretable and less aligned with the desiderata of job seekers. Departing from this approach, process fairness is operationalized with reference to the information on sensitive attributes encoded in the remaining variables. More work is required to understand process fairness in hiring more broadly and to suitably (if at all) operationalize it quantitatively. Impact fairness quantifies the downstream impacts of algorithmic outcomes on different populations; modeling important aspects of the surrounding socio-technical system, it goes beyond brittle algorithmic abstraction. Since these measures are exceedingly rare, we highlight the need for more context-specific research to understand and model the benefits and harms of decisions on job candidates at different stages of the hiring pipeline.
Blurred lines between ranking and classification. In the literature, classification and ranking are typically considered separate tasks with separate algorithmic fairness measures. In hiring, we find ranking systems evaluated according to classification measures such as $\operatorname{DI}$ (Burke et al., 2021), selection tasks solved with rankers (Burke et al., 2021), and target variables of similar nature encoded as binary, multinary, or continuous (Section 6). For this reason, the separation between classification and ranking is less rigid than in other domains, and the question of which measures are best suited to evaluate these systems arises. A fundamental part of the answer lies in the expected flexibility of use for these models. If the key aspects of their usage are fixed, including the cutoff thresholds for different outcomes, then a measure with coarse granularity, considering a single operating condition (typical of the fair classification literature), is most suited for evaluation. On the contrary, for models whose operating conditions are more uncertain, requiring human interaction or threshold fine-tuning, it is preferable to adopt a measure with finer granularity (often derived from the fair ranking literature), e.g. averaging outcomes across multiple realizations of a browsing model.
@@
Fairness diagnostics vs. fairness optimization. Ease of interpretation is a desirable property, especially for fairness measures, as confirmed by the popularity of DI. Nevertheless, several fairness measures adopted in hiring are difficult to read in absolute terms, making it challenging to understand the severity of fairness violations based on their values. These values can only be interpreted relative to one another, i.e. it is clear that one value is more desirable than another, making them a possible target for optimization but less suited for diagnostics. Measures of outcome and impact fairness tend to be more interpretable, since they are typically defined as differences between the probability of obtaining desirable outcomes for different groups. This is especially true for binary measures, while multinary measures, such as $\operatorname{RMS}$, often sacrifice interpretability to summarize disparities across multiple groups.
@@
Ignoring privileged and disadvantaged groups. The definition of sensitive attributes is informed by the historical (dis)advantage of specific groups. Gender and race, for example, are considered sensitive attributes due to the recurrent structural disadvantages incurred by women and black people. This is especially true in hiring, where surveys and meta-analyses find consistent biases against women and ethnic minorities (Bertrand and Mullainathan, 2004; Kline et al., 2022; Quillian et al., 2017). Although it is commonly held, especially in fairness research, that disparities against disadvantaged groups should be mitigated, there is less consensus on how to treat algorithms that happen to favor historically disadvantaged communities. In other words, should algorithmic fairness be symmetrical and reject a priori notions of privileged and disadvantaged groups? This is an important normative question, seldom acknowledged in the fairness literature, with immediate consequences on the choice of measures. For example, the version of $\operatorname{DI}$ reported in Equation (3) computes the ratio between the acceptance rates of the worse-off group ($\min$) and the best-off group ($\max$), ignoring prior knowledge of structural inequality and affected communities.
@@ 5.1. Pre-processing
in family measures
specifics*
approach summary
Rule-based Scraping or Substitution
(De-Arteaga et al., 2019; Parasurama and Sedoc, 2021; Rus et al., 2022)
PRE
$\operatorname{sAUC}$, $\operatorname{DD}$, $\operatorname{TPRD}$, $\operatorname{SD}$
text
proxy reduction
remove or substitute gender identifiers based on hard-coded rules
Importance-based Scraping
(Parasurama and Sedoc, 2021; Booth et al., 2021)
PRE
$\operatorname{sAUC}$, $\operatorname{DI}$
proxy reduction
iteratively remove proxy features most predictive of sensitive attribute
Balanced Sampling
(Arafan et al., 2022; Yan et al., 2020)
PRE
$\operatorname{NDKL}$, $\operatorname{MIA}$, $\operatorname{MAE}$
rebalancing
resample training set with same group cardinality in each class
Group Norming
(Booth et al., 2021)
PRE
$\operatorname{sAUC}$, $\operatorname{DI}$
groupwise feature transform
$z$-score normalization of each feature within each group
Subspace Projection
(Parasurama and Sedoc, 2021)
PRE
$\operatorname{sAUC}$
text
proxy reduction
remove gender information via embedding projection
Adversarial Inference
(Hemamou et al., 2021; Rus et al., 2022; Yan et al., 2020; Peña et al., 2020)
IN
$\operatorname{sAUC}$, $\operatorname{DD}$, $\operatorname{TPRD}$, $\operatorname{SD}$, $\operatorname{MIA}$, $\operatorname{MAE}$, $\operatorname{SKL}$
proxy reduction
reduce sensitive info in latent representation through additional adv. loss
Face Decorrelation
(Hemamou et al., 2021)
IN
$\operatorname{sAUC}$, $\operatorname{DI}$
image, $s$ unk.
proxy reduction
discourage intermediate representations correlated with face ID through adv. loss
Name Decorrelation
(Hemamou and Coleman, 2022)
IN
$\operatorname{sAUC}$, $\operatorname{TPRD}$, $\operatorname{RMS}$
text, $s$ unk.
proxy reduction
discourage intermediate representation correlated with names by minimizing MI
Fair TF-IDF
(Deshpande et al., 2020)
IN
$\operatorname{DI}$
text
proxy reduction
reduce feature weight according to its group-specificity
DetGreedy
(Geyik et al., 2019; Arafan et al., 2022; Sühr et al., 2021)
POST
$\operatorname{NDKL}$
output re-ranking
re-rank items enforcing desired group representation in each prefix
CDF Rescoring
(Nandy et al., 2022)
POST
$\operatorname{xEO}$
output re-rescoring
re-score items with groupwise CDF trick
Spatial Partitioning
(Burke et al., 2021)
POST
$\operatorname{DI}$
$s$ unk.
output re-ranking
promote candidates based on group membership probability
  • * method specifics indicate whether it is designed for a specific data modality (text or image) or a specific setting (sensitive attribute $s$ unknown)
@@ 5.4. Discussion
Outcome fairness through proxy reduction? A very clear trend in the literature is the popularity of proxy reduction methods that target process fairness ($\operatorname{sAUC}$) in pursuit of outcome fairness (e.g. $\operatorname{DI}$, $\operatorname{TPRD}$). This approach probably gained popularity because it aligns with legislation against disparate treatment (US) and direct discrimination (EU), which prohibit basing hiring decisions on protected attributes (Lambrecht and Tucker, 2019; Adams-Prassl et al., 2023) and because it does not require run-time access to sensitive attributes. This approach is also adopted by vendors, such as Pymetrics and Hirevue, upon detecting violations of the 80% rule in pre-deployment audits (Raghavan et al., 2020; HireVue, 2022). While this method can mitigate large outcome disparities, it does not produce convincing improvements when the inequity is less significant (Hemamou et al., 2021; Hemamou and Coleman, 2022; Booth et al., 2021). Moreover, it is worth noting that, in contrast to post-processing, this approach achieves mitigation on a held-out test set during development, but provides no guarantee at deployment. More general studies are required to understand the effects of proxy removal on outcome fairness and ensure actual benefits for vulnerable candidates.
@@
Opportunities and risks of post-processing. Post-processing approaches are the easiest to integrate into existing systems, as they can be modularly added to algorithmic pipelines (Vasudevan and Kenthapadi, 2020). Indeed, most post-processing methods in algorithmic hiring are proposed and deployed at LinkedIn, a large company with an established platform powered by interactions between complex data infrastructure and interdependent algorithmic modules (Geyik et al., 2018; Auradkar et al., 2012). It is worth noting that post-processing explicitly takes into account sensitive attributes to change algorithmic outcomes for job candidates, which may be critical for disparate treatment and affirmative action (US) as well as direct discrimination and positive action (EU) (Bent, 2019; Hacker, 2018). Additionally, postprocessing requires run-time access to the sensitive attributes of all data subjects. This are likely two key factors explaining why post-processing is less popular and why hybrid approaches combining different types of fairness interventions, e.g. mitigating via pre- and post-processing, remain under-explored (Arafan et al., 2022).

### v3

@@ Abstract.
copyright: rightsretainedjournal: TISTjournalyear: 2024journalvolume: 1journalnumber: 1article: 1publicationmonth: 1doi: 10.1145/3696457conference: Make sure to enter the correct conference title from your rights confirmation emai; September 2024; price: 15.00isbn: 978-1-4503-XXXX-X/18/06ccs: General and reference Surveys and overviewsccs: Information systems Data management systemsccs: Social and professional topics Computing / technology policy
{mybox}
Correct reference for this work:
Alessandro Fabris, Nina Baranowska, Matthew J. Dennis, David Graus, Philipp Hacker, Jorge Saldivar, Frederik Zuiderveen Borgesius, and Asia J. Biega. Fairness and Bias in Algorithmic Hiring: a Multidisciplinary Survey. ACM Transactions on Intelligent Systems and Technology. 2024.
@@ 4.2. Notation
$x\in\mathcal{X}$ a vector of non-sensitive attributes
$s\in\mathcal{S}$ a vector of sensitive attributes
$s=g$ a sensitive attribute value
$s={g}^{\mathsf{c}}$ complement of a sensitive attribute value
$y\in\mathcal{Y}$ target variable from domain $\mathcal{Y}$
$i=(x_{i},s_{i},y_{i})$ a data point or item
$\hat{y}=f(x)$ a classifier $f:\mathcal{X}\rightarrow\mathcal{Y}$ issuing predictions in $\mathcal{Y}$ for data points in $\mathcal{X}$
$f_{\text{soft}}(x)$ a soft classifier, from which predictions $\hat{y}=f(x)$ can be derived through thresholding
$\tau=\text{argsort}(f_{\text{soft}}(x))$ ranking of items, typically sorted by target estimates
$i=\tau(k)$ item at rank $k$ in $\tau$
$N$ number of items
$N_{g}$ number of items in $g$
$N_{g}^{k}$ number of items in $g$ in the top $k$ positions of a ranking
$D_{g}$ desired representation for group $g$
$h(x)$ a classifier $h:\mathcal{X}\rightarrow\mathcal{S}$ issuing predictions in $\mathcal{S}$ for data points
$h_{\text{soft}}(x)$ a soft classifier for sensitive attributes
@@ 4.3.2. Accuracy fairness
$\operatorname{NDKL}$ (Geyik et al., 2019; Arafan et al., 2022) outcome No
$\operatorname{DI}$ (Booth et al., 2021; Burke et al., 2021; Köchling et al., 2021, 2021; Wilson et al., 2021; Deshpande et al., 2020) outcome No
$\operatorname{DD}$ (Rus et al., 2022) outcome No No
$\operatorname{RPP}$ (Ali et al., 2019; Imana et al., 2021; Peña et al., 2020) outcome No No No
$\operatorname{DRD}$ (Zhang et al., 2022) outcome No
$\operatorname{SKL}$ (Peña et al., 2020) outcome No No
$\operatorname{LRR}$ (Chen et al., 2018a) outcome $x$ No No D
$\operatorname{TPRD}$ (De-Arteaga et al., 2019; Hemamou and Coleman, 2022) outcome $y$ No No
$\operatorname{FNRR}$ (Köchling et al., 2021) outcome $y$ No No
$\operatorname{MED}$ (Singhania et al., 2020) outcome $y$
$\operatorname{RMS}$ (Hemamou and Coleman, 2022) outcome $y$ No No No
$\operatorname{xEO}$ (Nandy et al., 2022) outcome $y$ No No No
$\operatorname{MAE}$ (Yan et al., 2020; Singhania et al., 2020) accuracy $y$ No No
$\operatorname{BCRD}$ (Köchling et al., 2021) accuracy $y$ No No No
$\operatorname{MID}$ (Köchling et al., 2021) accuracy $y$ No No No
$\operatorname{SD}$ (Rus et al., 2022) impact No No
$\operatorname{GBS}$ (Hu et al., 2022) representational No n.a No
$\operatorname{sAUC}$ (Parasurama et al., 2022; Parasurama and Sedoc, 2021; Booth et al., 2021; Hemamou et al., 2021; Rus et al., 2022; Hemamou and Coleman, 2022) process No n.a. No n.a.
$\operatorname{GTR}$ (Parasurama et al., 2022) process $x$ No n.a. No n.a.
$\operatorname{MIA}$ (Yan et al., 2020) process $y$ No n.a.
@@ 4.4. Discussion
In this section, we describe the main characteristics of the surveyed measures and how they inform fairness measurement choices.
[…]
To reiterate, this is just one fictitious scenario and not a fixed recommendation for choosing the right measure. It demonstrates the importance of key characteristics of the measures summarized in Table 2 for practitioner decisions; in the paragraphs below, we expand on the factors driving these decisions.
@@
Fairness diagnostics vs. fairness optimization. Ease of interpretation is a desirable property, especially for fairness measures, as confirmed by the popularity of DI. Nevertheless, several fairness measures adopted in hiring are difficult to read in absolute terms, making it challenging to understand the severity of fairness violations based on their values. These values can only be interpreted relative to one another, i.e. it is clear that one value is more desirable than another, making them a possible target for optimization but less suited for diagnostics. This is especially true for measures of accuracy fairness such as $\operatorname{BCRD}$ and $\operatorname{MID}$. Measures of outcome and impact fairness, such as $\operatorname{DD}$ and $\operatorname{SD}$, tend to be more interpretable, since they are typically defined as differences between the probability of obtaining desirable outcomes for different groups. This holds especially for binary measures, while multinary measures, such as $\operatorname{RMS}$, often sacrifice interpretability to summarize disparities across multiple groups.
Blurred lines between ranking and classification. In the literature, classification and ranking are typically considered separate tasks with separate algorithmic fairness measures. In hiring, we find ranking systems evaluated according to classification measures such as $\operatorname{DI}$ (Burke et al., 2021), selection tasks solved with rankers (Burke et al., 2021), and target variables of similar nature encoded as binary, multinary, or continuous (Section 6). For this reason, the separation between classification and ranking is less rigid than in other domains, and the question of which measures are best suited to evaluate these systems arises. A fundamental part of the answer lies in the expected flexibility of use for these models. If the key aspects of their usage are fixed, including the cutoff thresholds for different outcomes, then a measure with coarse granularity, considering a single operating condition (typical of the fair classification literature – e.g. $\operatorname{DD}$, $\operatorname{TPRD}$), is most suited for evaluation. On the contrary, for models whose operating conditions are more uncertain, requiring human interaction or threshold fine-tuning, it is preferable to adopt a measure with finer granularity (often derived from the fair ranking literature – e.g. $\operatorname{NDKL}$, $\operatorname{DRD}$), which averages outcomes across multiple realizations of a browsing model.
Diagnosing biases. Fairness measures are especially useful when they provide actionable diagnostics by highlighting specific biases described in Section 3. Measures of process fairness such as $\operatorname{sAUC}$ and $\operatorname{MIA}$ detect the presence of sensitive attribute proxies. Dedicated proxy reduction approaches (Section 5) can mitigate the resulting risk of discrimination. Representational fairness ($\operatorname{GBS}$) captures biases in representations; when applied to job descriptions, extreme values highlight stereotypically gendered language, signaling a risk of culture-based avoidance. Impact fairness can highlight nuanced aspects of hiring beyond selection rates; $\operatorname{SD}$, for instance, captures wage gaps and salary negotiation differences prompting further analysis into remuneration packages. Measures of outcome (e.g. $\operatorname{DI}$) and accuracy fairness (e.g. $\operatorname{MAE}$) can serve as high-level metrics by highlighting the effects of multiple factors acting together. To exemplify, a combination of job segregation, work gaps, and disparate performance of language processing tools, even if relatively weak on their own, may trigger a very low (unfair) value of DI when they compound.
Assorted fairness flavors. No single measure provides a complete picture. Different fairness flavors have complementary strengths and weaknesses. Most of the fairness measures considered in hiring focus on outcome equity; the field is strongly influenced by the 80% rule (Section 8.2), which is appreciated as a quantitative rule of thumb, but also often criticized (Wax, 2011; EEOC - US Equal Employment Opportunity Commission, 2023). Indeed, outcome fairness is based on a narrow view of algorithmic hiring as a single decision point abstracted away from its context, where equity is purely a function of algorithmic estimates and (sometimes) their similarity to target variables approximating a ground truth. Similar criticism can be moved to accuracy fairness, with the additional limitation that it is less interpretable and less aligned with the desiderata of job seekers. Departing from this approach, process fairness is operationalized with reference to the information on sensitive attributes encoded in the remaining variables. More work is required to understand process fairness in hiring more broadly and to suitably (if at all) operationalize it quantitatively. Impact fairness quantifies the downstream impacts of algorithmic outcomes on different populations; modeling important aspects of the surrounding socio-technical system, it goes beyond brittle algorithmic abstraction. Since these measures are exceedingly rare, we highlight the need for more context-specific research to understand and model the benefits and harms of decisions on job candidates at different stages of the hiring pipeline. Overall, practitioners developing equity monitoring protocols should consider multiple fairness angles to gain a nuanced and more complete picture.
@@
Ignoring privileged and disadvantaged groups. The definition of sensitive attributes in anti-discrimination law is informed by the historical (dis)advantage of specific groups (Simson et al., 2024). Gender and race, for example, are considered sensitive attributes due to the recurrent structural disadvantages incurred by women and black people. This is especially true in hiring, where surveys and meta-analyses find consistent biases against women and ethnic minorities (Bertrand and Mullainathan, 2004; Kline et al., 2022; Quillian et al., 2017). Although it is commonly held, especially in fairness research, that disparities against disadvantaged groups should be mitigated, there is less consensus on how to treat algorithms that happen to favor historically disadvantaged communities. In other words, should algorithmic fairness be symmetrical and reject a priori notions of privileged and disadvantaged groups? This is an important normative question, seldom acknowledged in the fairness literature, with immediate consequences on the choice of measures. For example, the version of $\operatorname{DI}$ reported in Equation (3) computes the ratio between the acceptance rates of the worse-off group ($\min$) and the best-off group ($\max$), ignoring prior knowledge of structural inequality and affected communities.
@@ 5.1. Pre-processing
in family measures $s$ availability mode approach summary
Rule-based Scraping or Substitution (De-Arteaga et al., 2019; Parasurama and Sedoc, 2021; Rus et al., 2022) PRE $\operatorname{sAUC}$, $\operatorname{DD}$, $\operatorname{TPRD}$, $\operatorname{SD}$ none text proxy reduction remove or substitute gender identifiers based on hard-coded rules
Importance-based Scraping (Parasurama and Sedoc, 2021; Booth et al., 2021) PRE $\operatorname{sAUC}$, $\operatorname{DI}$ train proxy reduction iteratively remove proxy features most predictive of sensitive attribute
Balanced Sampling (Arafan et al., 2022; Yan et al., 2020) PRE $\operatorname{NDKL}$, $\operatorname{MIA}$, $\operatorname{MAE}$ train rebalancing resample training set with same group cardinality in each class
Group Norming (Booth et al., 2021) PRE $\operatorname{sAUC}$, $\operatorname{DI}$ train groupwise feature transform $z$-score normalization of each feature within each group
Subspace Projection (Parasurama and Sedoc, 2021) PRE $\operatorname{sAUC}$ none text proxy reduction remove gender information via embedding projection
Adversarial Inference (Hemamou et al., 2021; Rus et al., 2022; Yan et al., 2020; Peña et al., 2020) IN $\operatorname{sAUC}$, $\operatorname{DD}$, $\operatorname{TPRD}$, $\operatorname{SD}$, $\operatorname{MIA}$, $\operatorname{MAE}$, $\operatorname{SKL}$ train proxy reduction reduce sensitive info in latent representation through additional adv. loss
Face Decorrelation (Hemamou et al., 2021) IN $\operatorname{sAUC}$, $\operatorname{DI}$ none image proxy reduction discourage intermediate representations correlated with face ID through adv. loss
Name Decorrelation (Hemamou and Coleman, 2022) IN $\operatorname{sAUC}$, $\operatorname{TPRD}$, $\operatorname{RMS}$ none text proxy reduction discourage intermediate representation correlated with names by minimizing MI
Fair TF-IDF (Deshpande et al., 2020) IN $\operatorname{DI}$ train text proxy reduction reduce feature weight according to its group-specificity
DetGreedy (Geyik et al., 2019; Arafan et al., 2022; Sühr et al., 2021) POST $\operatorname{NDKL}$ runtime output re-ranking re-rank items enforcing desired group representation in each prefix
CDF Rescoring (Nandy et al., 2022) POST $\operatorname{xEO}$ runtime output re-rescoring re-score items with groupwise CDF trick
Spatial Partitioning (Burke et al., 2021) POST $\operatorname{DI}$ train output re-ranking promote candidates based on group membership probability
@@ 5.4. Discussion
In this section, we present the main trends for fairness enhancement in the algorithmic hiring literature and describe important factors guiding these choices (summarized in Table 3). Section 7 will expand on this analysis presenting additional dimensions that guide fairness mitigation in practice.
Mitigating biases. Some of the proposed approaches are suited to mitigate specific biases described in Section 3. Balanced sampling, for example, can counter the effect of under-representation in training sets caused by factors such as horizontal and vertical job segregation. Proxy reduction methods (e.g. decorrelation, adversarial inference) target sensitive attribute proxies, reducing the overreliance of models on sensitive information. Group norming seems applicable to counter systematic biases encoded in input features against protected groups, such as biased employee evaluations. However, it is certainly preferable to understand and improve the measurement system that provides these biased features rather than simply matching the distribution of input features across protected groups without a clear understanding of the underlying socio-technical system. Finally, output re-ranking can mitigate the joint effect of different bias conducive factors. Its application comes with risks and opportunities described below.
@@
Opportunities and risks of post-processing. Post-processing approaches, such as DetGreedy, are the easiest to integrate into existing systems, as they can be modularly added to algorithmic pipelines (Vasudevan and Kenthapadi, 2020). Indeed, most post-processing methods in algorithmic hiring are proposed and deployed at LinkedIn, a large company with an established platform powered by interactions between complex data infrastructure and interdependent algorithmic modules (Geyik et al., 2018; Auradkar et al., 2012). It is worth noting that post-processing explicitly takes into account sensitive attributes to change algorithmic outcomes for job candidates, which may be critical for disparate treatment and affirmative action (US) as well as direct discrimination and positive action (EU) (Bent, 2019; Hacker, 2018). Additionally, postprocessing requires run-time access to the sensitive attributes of all data subjects, as described in the next paragraph. These are likely two key factors explaining why post-processing is less popular and why hybrid approaches combining different types of fairness interventions, e.g. mitigating via pre- and post-processing, remain under-explored (Arafan et al., 2022).
Sensitive data. Different mitigation strategies have different requirements for sensitive attributes $s$. Post-processing methods, such as DetGreedy and CDF restoring, typically require knowledge of sensitive attributes during runtime. Conversely, pre-processing and in-processing approaches, such as importance-based scraping and adversarial inference, require access to sensitive attributes only during training, not testing. Furthermore, specific data modalities, such as text and vision, facilitate specialized methods like subspace projection and face decorrelation, which function without sensitive attribute information about data subjects. Each of these methods has distinct data requirements, ranging from highly restrictive to more lenient. Having access to sensitive attributes at runtime enables precise and reliable interventions, whereas strategies that do not rely on this data must be continuously monitored to ensure they are meeting their intended goals.
Outcome fairness through proxy reduction? A very clear trend in the literature is the popularity of proxy reduction methods, such as adversarial inference, that target process fairness ($\operatorname{sAUC}$) in pursuit of outcome fairness (e.g. $\operatorname{DI}$, $\operatorname{TPRD}$). This approach probably gained popularity because it aligns with legislation against disparate treatment (US) and direct discrimination (EU), which prohibit basing hiring decisions on protected attributes (Lambrecht and Tucker, 2019; Adams-Prassl et al., 2023) and because it does not require run-time access to sensitive attributes. This approach is also adopted by vendors, such as Pymetrics and Hirevue, upon detecting violations of the 80% rule in pre-deployment audits (Raghavan et al., 2020; HireVue, 2022). While this method can mitigate large outcome disparities, it does not produce convincing improvements when the inequity is less significant (Hemamou et al., 2021; Hemamou and Coleman, 2022; Booth et al., 2021). Moreover, it is worth noting that, in contrast to post-processing, this approach achieves mitigation on a held-out test set during development, but provides no guarantee at deployment. More general studies are required to understand the effects of proxy removal on outcome fairness and ensure actual benefits for vulnerable candidates.